find the equation of each line which passes through the point (3,2) and is parallel to y = 3x - 5. how do i solve this??
Answer by checkley75(3666)

Y=mX+b WHERE m=SLOPE & b=Y INTERCEPT.
Y=3X-5 this line has a slope (m) of 3.
TO FIND THE EUATION OF THE LINE THROUGH (3,2) WITH A SLOPE OF 3 WE REPLACE THE X & Y VALUES WITH 3 & 2 TO FIND THE Y INTERCEPT (b) THUS:
2=3*3+b
2=9+b
-7=b THEN THE EQUATION OF THIS LINE IS:
Y=3X-7